The position of a certain electron is determined at a certain time with an uncertainty of
10-10 m. Find the uncertainty in the electron's momentum. (2 marks)
1
Expert's answer
2020-06-30T18:23:51-0400

According to the uncertainty principle, "\sigma_x\sigma_p \ge \dfrac{\hbar}{2}." Therefore, "\sigma_p \ge \dfrac{\hbar}{2} \cdot\dfrac{1}{\sigma_x} = \dfrac{1.05\cdot10^{-34}}{2}\cdot\dfrac{1}{10^{-10}} = 5.25\cdot10^{-25}\,\mathrm{kg\, m/s}."